# Dissipation alters modes of information encoding in small quantum reservoirs near criticality

### v1

@@ 1 Motivation and Introduction
Reservoir Computing (RC) is a computational paradigm that harnesses the intrinsic dynamics of complex systems to process time-dependent inputs efficiently [1, 2, 3]. Unlike conventional recurrent neural networks (RNNs), RC requires training only at the readout layer, circumventing expensive weight-update procedures on internal nodes [4]. Quantum Reservoir Computing (QRC) extends these ideas to quantum platforms, leveraging quantum superposition and entanglement to amplify the dimensionality of the feature space and potentially enhance computational capabilities [5, 3]. Early demonstrations of QRC have shown promise in tasks like time-series prediction, classification, and memory capacity estimation, and ongoing efforts explore a range of theoretical and experimental strategies for improving performance [6, 7, 8, 9, 10, 11, 12, 13, 14].
@@
To study the information encoding properties of this system, we borrow information theoretic tools from neuroscience, where the concepts of synergy and redundancy of information encoding have been used to understand how groups of neurons encode stimuli [18, 19, 20, 21]. Redundancy measures the overlap of information shared among components, while synergy quantifies information that emerges only when components are considered together. These ideas are particularly relevant for QRC, where the interaction with quantum systems determines how information is processed. However, traditional information measures like mutual information cannot disentangle synergy and redundancy. To address this, we employ Partial Information Decomposition (PID) [22], which decomposes mutual information into non-overlapping components of synergistic, redundant, and unique information.
@@
Our main findings reveal that near a critical coupling strength leading to dynamical bifurcation, the system transitions from predominantly redundant encoding to a regime featuring significant synergistic information. This synergistic behavior arises from the interplay between fast collective oscillations and overdamped soft modes. We show that increasing dissipation, while suppressing quantum correlations, can nevertheless promotes highly redundant encoding modes that improve long-term memory retention. In contrast, near the onset of dynamical instability, synergy is amplified and enriches short-term responsiveness, improving short-term memory retention. Taken together, these results highlight how dissipation and dynamic instability in a minimal system can steer a quantum reservoir toward redundant or synergistic processing, each regime benefiting different computational tasks.
@@ 2.3 Characterizing Information Processing
The PID analysis focuses on classical correlations of input-output classical variables, but our system is fundamentally quantum. To probe quantum correlations inherent in our system, we measure the quantum mutual information (QMI)
$I(1:2)=S_{1}+S_{2}-S_{12},$ (7)
where $S_{i}$ is the von Neumann entropy of the reduced density matrix $\hat{\rho}_{i}=\mathrm{Tr}_{j\neq i}(\hat{\rho}_{12})$ for the $i^{\text{th}}$ oscillator, and $S_{12}$ is the von Neumann entropy of the joint two-oscillator state $\hat{\rho}_{12}$.
To compute these entropies, we first obtain $\hat{\rho}_{12}$ from the numerical simulation, then trace out one oscillator to get the reduced density matrix $\hat{\rho}_{i}$. Since the von Neumann entropy of a density matrix $\hat{\rho}$ is defined as $S(\hat{\rho})=-\mathrm{Tr}(\hat{\rho}\log\hat{\rho})=-\sum_{k}\lambda_{k}\log\lambda_{k},$ where $\{\lambda_{k}\}$ are the eigenvalues of $\hat{\rho}$, we first extract eigenvalue spectra of $\hat{\rho}_{12}$, $\hat{\rho}_{1}$ and $\hat{\rho}_{2}$ and then calculate $S_{1}$, $S_{2}$, and $S_{12}$. The resulting QMI indicates the degree of quantum correlations between the two quantum oscillators.
Memory Capacity (MC).
In addition to instantaneous encoding, we are interested in how the system stores information over time, as is relevant in quantum reservoir computing (QRC). The memory capacity quantifies how well the current state of the reservoir (the outputs $X_{1}(t)$, $X_{2}(t)$) retains information about past inputs $s(t-\tau)$. By analyzing how the reconstruction error of past inputs varies with the delay $\tau$, we derive a memory measure that complements the PID and QMI analyses.
A high memory capacity suggests that the reservoir not only encodes the input at a given instant but also preserves information over extended periods. Comparing memory capacity with $\mathrm{Rdn}$ and $\mathrm{Syn}$ reveals how different dynamical regimes influence both the instantaneous and temporal dimensions of information processing in this system. Details of MC calculation is provided in Sec. 3.4.
@@ 3.2 Underlying Mechanisms Enhancing Synergistic Coding: The Role of Soft and Fast Modes
Figure 4: Effective potential around the steady state $\alpha_{1,c}=\alpha_{2,c}=0$ (red dot) in the $\Delta<0,J>0$ regime, projected onto $\mathrm{Im}(\alpha_{1})=\mathrm{Im}(\alpha_{2})=0$. As $J$ crosses $|\Delta|$, flat directions appear, signaling soft or marginal modes.
@@
Figure 5: Retarded Green’s function poles, plotted as $J$ increases (Eq. (31)). Orange dots represent slow modes $\omega_{s}$, and blue dots represent fast modes $\omega_{f}$. The real part of $\omega_{s}$ approaches zero near $J=|\Delta|$, signaling the disappearance of slow collective oscillations and an overdamped decay to the steady state. The whole relaxation dynamics is approximately an underdamped dynamics dominated by the fast oscillatory modes.
@@ 3.4 Memory Capacity of Synergistic and Redundant Encoding
At higher dissipation rates ($\gamma$), the system shifts toward redundant encoding. In this regime, the reservoir rapidly relaxes to a steady state, which suppresses short-term memory performance. However, redundancy improves long-term memory capacity as information is encoded near the steady-state configuration. This shift reflects dissipation’s dual role: it acts both as a stabilizing force that ensures encoding near the steady state and as a driver of redundancy, which favors longer-term information retention in the simple uniformly random input memorization task.
From the perspective of quantum reservoir design, dissipation plays a critical role in ensuring the reservoir exhibits the fading memory property, where the influence of past inputs gradually diminishes, which is a necessary condition for designing operational reservoir computing [10, 9].
@@ 4 Conclusion and Outlook
Finally, it is important to recognize that the results presented here, like much of the prior work, assume perfect readout without the presence of shot noise. As system sizes scale, however, readout processes may suffer from exponential concentration phenomena, requiring exponentially many measurement shots to estimate input-dependent readouts accurately, as discussed in [11]. This limitation presents a significant barrier to scalability. In larger quantum reservoir systems, future studies must address how synergy and redundancy behave under realistic measurement constraints. Incorporating the effects of finite measurement precision into the framework of encoding dynamics could lead to more scalable and experimentally feasible designs for quantum reservoirs.

### v2

@@ 1 Motivation and Introduction
Reservoir Computing (RC) is a computational paradigm that harnesses the intrinsic dynamics of complex systems to process time-dependent inputs efficiently [1, 2, 3]. Unlike conventional recurrent neural networks (RNNs), RC requires training only at the readout layer, circumventing expensive weight-update procedures on internal nodes [4]. Quantum Reservoir Computing (QRC) extends these ideas to quantum platforms, leveraging quantum superposition and entanglement to amplify the dimensionality of the feature space and potentially enhance computational capabilities [5, 3]. Early demonstrations of QRC have shown promise in tasks like time-series prediction, classification, and memory capacity estimation, and ongoing efforts explore a range of theoretical and experimental strategies for improving performance [6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
@@
Our main findings reveal that near a critical coupling strength leading to dynamical bifurcation, the system transitions from predominantly redundant encoding to a regime featuring significant synergistic information. This synergistic behavior arises from the interplay between fast collective oscillations and overdamped soft modes. We show that increasing dissipation suppresses quantum correlations, and promotes highly redundant encoding modes. In contrast, near the onset of dynamical instability, synergy is amplified and enriches short-term responsiveness, improving short-term memory retention. Taken together, these results highlight how dissipation and dynamic instability in a minimal system can steer a quantum reservoir toward redundant or synergistic processing, each regime benefiting different computational tasks.
In relation to recent proposals using two coupled Kerr-nonlinear oscillators as quantum reservoirs [25, 26], which highlight the roles of dissipation for fading memory and moderate coupling for richer dynamics, our work differs by focusing on PID to examine how synergy and redundancy influences the reservoir’s memory capacity. This complements prior findings to show that critical points in Kerr dynamics can shift encoding from redundant to synergistic regimes. Meanwhile, single Kerr oscillators with large Hilbert spaces [13, 27] highlight how dimensionality alone can serve as a computational resource, but our key question of whether the whole can exceed the sum of its parts requires at least two coupled oscillators for emergent synergistic encoding. Lastly, although [28] studies larger arrays of Kerr oscillators and demonstrates near-bifurcation enhancements of nonlinear memory and employ higher-order cumulant expansions to handle higher photon numbers, we limit ourselves to at most a second-order expansion in a parameter regime where it remains accurate, focusing on PID-based insights rather than reservoir benchmarks and thus complementing other findings in the literature.
@@ 2.3 Characterizing Information Processing
Although our primary partial information decomposition (PID) analysis focuses on classical correlations between input-output variables, the underlying reservoir dynamics remain fundamentally quantum. To probe quantum correlations inherent in our system, we compute the quantum mutual information (QMI) between the two oscillators.
Let $\rho_{12}$ be the density matrix describing the joint state of the two coupled Kerr oscillators. We partition the system into subsystems $1$ and $2$, each corresponding to one oscillator. The QMI is then given by
$I(1:2)\;=\;S(\rho_{1})\;+\;S(\rho_{2})\;-\;S(\rho_{12}),$ (7)
where $\rho_{1}=\mathrm{Tr}_{2}(\rho_{12})$ and $\rho_{2}=\mathrm{Tr}_{1}(\rho_{12})$ are the reduced density matrices of each oscillator, and $S(\rho)=-\mathrm{Tr}\!\bigl{(}\rho\ln\rho\bigr{)}$ is the von Neumann entropy.
Because each oscillator’s Hilbert space is, in principle, infinite-dimensional, we employ a finite photon-number basis up to a cutoff $N_{\mathrm{cutoff}}$ to ensure computational tractability. Namely,
$\rho_{12}\;\approx\;\sum_{n_{1}=0}^{N_{\mathrm{cutoff}}}\sum_{n_{2}=0}^{N_{\mathrm{cutoff}}}c_{n_{1},n_{2}}\,|n_{1}\rangle\langle n_{1}|\;\otimes\;|n_{2}\rangle\langle n_{2}|,$ (8)
where $|n_{i}\rangle$ is the Fock state with $n_{i}$ photons in oscillator $i$, and $c_{n_{1},n_{2}}$ are the matrix elements obtained by time-averaging the steady-state solution of the master equation. We verify that increasing $N_{\mathrm{cutoff}}$ beyond $\sim 10$ does not significantly alter the results within the quantum regime studied in this work, suggesting that the computed QMI converges to within numerical precision.
To evaluate Eq. (7), we first diagonalize the full two-oscillator density matrix $\rho_{12}$
$\rho_{12}\;=\;\sum_{k}\lambda_{k}\,\bigl{|}\psi_{k}\bigr{\rangle}\bigl{\langle}\psi_{k}\bigr{|},\quad S(\rho_{12})\;=\;-\sum_{k}\lambda_{k}\,\ln\lambda_{k},$ (9)
where $\{\lambda_{k}\}$ and $\{\lvert\psi_{k}\rangle\}$ are the eigenvalues and eigenbases of $\rho_{12}$, respectively. We then obtain the reduced density matrices for each oscillator by tracing out the other as $\rho_{1}=\mathrm{Tr}_{2}\bigl{(}\rho_{12}\bigr{)},$ and $\rho_{2}=\mathrm{Tr}_{1}\bigl{(}\rho_{12}\bigr{)}.$ Both $\rho_{1}$ and $\rho_{2}$ are similarly diagonalized to compute their von Neumann entropies, $S(\rho_{1})$ and $S(\rho_{2})$. Substituting these into Eq. (7) gives the QMI. The resulting $I(1:2)$ quantifies the total correlations between the two oscillators, including both classical and quantum components, and thus complements a classical PID-based analysis.
@@ 3.2 Underlying Mechanisms Enhancing Synergistic Coding: The Role of Soft and Fast Modes
Figure 4: Effective potential around the steady state $\alpha_{1,c}=\alpha_{2,c}=0$ (red dot) in the $\Delta<0,J>0$ regime, projected onto $\mathrm{Im}(\alpha_{1})=\mathrm{Im}(\alpha_{2})=0$. (Left) When $J<|\Delta|$, the steady state is weakly unstable in all directions, with no soft modes present, and the system predominantly encodes inputs redundantly. (Center) At the critical point $J=|\Delta|$, flat directions appear, marking the onset of soft modes. In this near-critical regime, collective oscillations enhance synergistic encoding. (Right) For $J>|\Delta|$, the potential deforms into a saddle, with two stable and two unstable directions. Here, the soft modes again disappear, and the system encodes inputs redundantly.
@@
Figure 5: Retarded Green’s function poles (33) in the complex-frequency plane as $J$ increases, illustrating the evolution of slow modes $\omega_{s}$ (orange dots) and fast modes $\omega_{f}$ (blue dots). For $J<|\Delta|$, both slow and fast modes coexist, and the system tends to encode inputs more redundantly. Near the critical point $J=|\Delta|$, the real part of $\omega_{s}$ approaches zero, indicating a disappearance of slow collective oscillations and an overdamped decay. In this near-critical regime, collective oscillations due to underdamped fast modes dominate and enhance synergistic encoding. For $J>|\Delta|$, the slow modes shift away from zero frequency, reducing synergy and transitioning the system back toward redundant encoding. This interplay between slow and fast modes governs how the system transitions from predominantly redundant to synergistic processing and back again as $J$ crosses the critical point.
@@ 3.4 Memory Capacity of Synergistic and Redundant Encoding
We note that, for low dissipation (e.g., $\gamma=0.5$), the system may appear to retain information about initial conditions at long delays $n$, indicating that the fading memory property might not be fully realized. While a more comprehensive benchmark such as the information processing capacity (IPC) [41] could more definitively confirm or refute fading memory in this low-dissipation regime, our current goal is to investigate synergy and redundancy in a small quantum reservoir rather than to perform an exhaustive reservoir-computing benchmark. Instead, we emphasize qualitative trends in short- vs. long-term memory. A more detailed, IPC-based study would be an interesting direction for future work to fully study the computational potential of near-critical quantum reservoirs in this system.
@@
At higher dissipation rates $\gamma$, the system tends to encode inputs more redundantly and relaxes more quickly to a steady state. Interestingly, as illustrated in Figure 9, both short- and long-term memory capacities, at criticality, increase with $\gamma$ for a simple uniformly random input memorization task, even though one might initially expect a trade-off. The underlying reason could be that the system has not yet fully achieved the fading memory regime, so further investigations, potentially through more comprehensive benchmarks such as IPC analysis, are needed to confirm how dissipation precisely balances short-term responsiveness and long-term stability right at a critical point.
Also from the perspective of quantum reservoir design, dissipation plays a critical role in ensuring the reservoir exhibits the fading memory property, where the influence of past inputs gradually diminishes, which is a necessary condition for designing operational reservoir computing [10, 9].
@@ 4 Conclusion and Outlook
In practice, realistic quantum devices inevitably face various noise sources and control imperfections. While these perturbations may shift the crossover point between redundant and synergistic encoding or relocate the system’s critical point, our results show that this crossover in the modes of encoding persists over a wide parameter range. With advanced experimental techniques, such as dynamical decoupling and error mitigation in quantum optics, we expect that the essential physics behind this crossover can still be realized in real experimental systems.
Lastly, it is important to recognize that the results presented here, like much of the prior work, assume perfect readout without the presence of shot noise. As system sizes scale, however, readout processes may suffer from exponential concentration phenomena, requiring exponentially many measurement shots to estimate input-dependent readouts accurately, as discussed in [11]. This limitation presents a significant barrier to scalability. In larger quantum reservoir systems, future studies must address how synergy and redundancy behave under realistic measurement constraints. Incorporating the effects of finite measurement precision into the framework of encoding dynamics could lead to more scalable and experimentally feasible designs for quantum reservoirs.